OpenStudy (anonymous):

Anyone up for this, 3x-y=12 x+3y=14 , solve by substitution.

OpenStudy (anonymous):

first we find the value of y from eq. I -y=12-3x so y=3x-12 now put this value in eq II... on putting we get x+3(3x-12)=14 = x+9x-36=14 = 10x-36=14 = 10x=14+36 = 10x=50 = x=5 now put x=5 in eq I then we get 3x-y=12 = 3*5-y=12 = 15-y=12 =y= 15-12 =y=3 finally x=5 & y=3

OpenStudy (radar):

Definitely rates a good answer, am awarding a medal for this step by step lucid solution.

OpenStudy (anonymous):

thanx dear i will be happy to solve more questions...

OpenStudy (anonymous):

Thank you